Circular No. 7575

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001K IN IC 677 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001K (cf. IAUC 7565), obtained by P. Berlind on Jan. 24.48 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova near maximum light. The spectrum consists of a blue continuum with superposed P-Cyg lines of hydrogen and Ca II. Adopting the NASA/IPAC Extragalactic Database recession velocity of 3249 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 8500 km/s.

XTE J1550–564 J. A. Tomsick, University of California at San Diego (UCSD); E. Smith

and J. Swank, Goddard Space Flight Center; R. Wijnands, Massachusetts Institute of Technology; and J. Homan, University of Amsterdam, report: “Recent observations of the blackhole x-ray transient XTE J1550–564 with the Rossi X-ray Timing Explorer (RXTE) indicate that the source is active again. The last outburst ended in 2000 August, and a Chandra observation made on 2000 Aug. 16 indicated a flux level of 1 × 10^{−13} erg cm^{−2} s^{−1} (0.5–8 keV). A 2.5–20-keV source flux of 9 × 10^{−10} erg cm^{−2} s^{−1} was measured by RXTE on 2001 Jan. 28.7 UT. The spectral and timing properties indicate that the source is in the low-hard state. The energy spectrum can be described by a power law with a photon index of 1.52 and a neutral iron line with an equivalent width of 160 eV. The level of timing noise is ∼ 40-percent rms (0.01–100 Hz; 2–60 keV), with violent flaring activity having a repetition period of 100–160 s. Currently, it is not clear if the observed activity will lead to a full outburst. Additional RXTE observations are scheduled, and coordinated observations at other wavelengths are encouraged.”

R. Jain and C. Bailyn, Yale University; and J. Tomsick, UCSD, report YALO consortium observations of the optical counterpart of XTE J1550– 564 using the Yale 1-m telescope (+ ANDICAM instrument) at Cerro Tololo Interamerican Observatory: “XTE J1550–564 has entered a new optical outburst. By Jan. 28.34, 29.34, and 30.34 UT, the object was at V = 18.28 ± 0.01, 18.27 ± 0.01, and 18.25 ± 0.01, respectively. The source is significantly brighter than the quiescent level of V ∼ 21.6.”

2001 January 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green